OpenStudy (anonymous):

Can someone solve (x+h)^3. Please show the step-by-step.

OpenStudy (anonymous):

(x+h)^3 = (x+h)(x+h)(x+h) = (x+h)[x^2 + 2xh + h^2) = x^3 + 2x^2h + xh^2 + hx^2 + 2xh^2 + h^3 = x^3 + 3x^2h + 3xh^2 + h^3